40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 142 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  9924   Wed May 7 22:47:33 2014 ranaUpdateCDScdsutils updated to version 226: not working on pianosa or rossa

 controls@rossa:~ 0$ cdsutils read C1:LSC-DARM_GAIN
Traceback (most recent call last):
  File "/usr/lib/python2.6/runpy.py", line 122, in _run_module_as_main
    "__main__", fname, loader, pkg_name)
  File "/usr/lib/python2.6/runpy.py", line 34, in _run_code
    exec code in run_globals
  File "/ligo/apps/cdsutils-226/lib/python2.6/site-packages/cdsutils/__main__.py", line 57, in <module>
    imp.load_module('__main__', f, pathname, description)
  File "/ligo/apps/cdsutils-226/lib/python2.6/site-packages/cdsutils/read.py", line 32, in <module>
    print ezca.Ezca(prefix).read(rest, as_string=args.as_string)
  File "/ligo/apps/linux-x86_64/cdsutils-226/lib/python2.6/site-packages/ezca/cached.py", line 17, in __call__
    key = (args, tuple(kwargs.viewitems()))
AttributeError: 'dict' object has no attribute 'viewitems'

  9923   Wed May 7 17:10:59 2014 ranaUpdateCDScdsutils updated to version 226

 This upgrade from Jamie has given us the new apps (avg, servo, and trigservo). We should figure out if there's a way to integrate Masayuki's work, so that we can have a 'cdsutils demod' function too.

  9922   Wed May 7 16:31:12 2014 jamieUpdateCDScdsutils updated to version 226
controls@pianosa:~ 0$ cd /opt/rtcds/cdsutils/trunk/
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ svn update
...
At revision 226.
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ make
echo "__version__ = '226'" >lib/cdsutils/_version.py
echo "__version__ = '226'" >lib/ezca/_version.py
...
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ make ligo-install
python ./setup.py install --prefix=/ligo/apps/linux-x86_64/cdsutils-226
...
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ ln -sfn cdsutils-226 /ligo/apps/linux-x86_64/cdsutils
controls@pianosa:/opt/rtcds/cdsutils/trunk 0$ exit
...
controls@pianosa:~ 0$ cdsutils --version
cdsutils 226
controls@pianosa:~ 0$ 

  9921   Wed May 7 14:01:36 2014 steveUpdateLSCTRY 60Hz noise hunt

Quote:

This is an effort to get rid of our ground loops  by isolating the electronic components from the optical table.

Aluminum mounting base plates of Thorlabs BA2 and Newport B-2 were replaced by plates or post made out of delrin material.

This is an insulator. DELRIN base plates were installed 6 places. The oplev-qpd has Nylon base plate.

The NPRO and HE/NE lasers are not isolated from the table. S8 and S9

I'm not sure about the doubling oven S10 

The optical table is grounded at G11  through  ~1 Mohms to the ETMY chamber.

Alignment touch up needed   at all D-marked component!

 

 Attachment appendix:

 

D: component delrin isolated

N: component nylon isolated ( or Delrin )

S: component shell is shorting to optical table (except oven)

G:  optical table ground

 

I failed to maximize TRY the pds.

  9920   Wed May 7 04:01:44 2014 rana, jenneUpdateLSCPRFPMI: Common Mode servo using REFL_DC ON, CARM offset still non-zero
  1. With REFL_DC coupled into the CM board through an SR560 (with an offset subtractor), we were able to transition to use it as the CARM error signal.
  2. We reduced the CARM offset until the arm powers went up to ~13.
  3. We had the AO path turned on and the MCL/AO crossover was ~150 Hz.
  4. We saw the double cavity pole come in from HF down to ~1-2 kHz. The lock stayed stable like this.
  5. We've set the IMC overall gain higher by +4dB in the mcup script. That's -4 dB from Eric's max gain earlier today.
  6. We have some scripts now for this scripts/PRFPMI/ :   camr_cm_down.sh and carm_cm_up.sh
  7. The sequence was ALS -> SqrtInv while digital with CARM -> MC2. Then we digital transition to REFL_DC using the CM board switch to put REFL_DC into the REFL11_I socket.
  8. REFL_DC is noisy, so we upped the SR560 gain by 10 and compensated.

Also, we found the PRM OL off and turned it back on. The ETMY was swinging a lot after lock loss, so we set its SUSPOS damping gain to match the ETMX and it stopped swinging so much.

Next up: more of the same, make this sequence more stable, turn on CARM OSC and watch the LOCKI outputs while we slowly ramp between signals.

Also, what should be the sign of the CARM offset ???

  9919   Tue May 6 19:38:13 2014 JenneUpdateLSCSet up for PRFPMI CM locking

To get ready for the PRFPMI CM trials tonight, I put AS55's cables back to their nominal state, and now have REFL11 I going to IN1 of the CM board.  The OUT1 of the CM board goes to the REFL11I whitening channel.

REFLDC was not disconnected in the last few days, so it is still set up for IN2 of the CM board, with an external offset adjust.

  9918   Tue May 6 18:32:14 2014 steveUpdateLSCTRY 60Hz noise hunt

This is an effort to get rid of our ground loops  by isolating the electronic components from the optical table.

Aluminum mounting base plates of Thorlabs BA2 and Newport B-2 were replaced by plates or post made out of delrin material.

This is an insulator. DELRIN base plates were installed 6 places. The oplev-qpd has Nylon base plate.

The NPRO and HE/NE lasers are not isolated from the table. S8 and S9

I'm not sure about the doubling oven S10 

The optical table is grounded at G11  through  ~1 Mohms to the ETMY chamber.

Alignment touch up needed   at all D-marked component!

 

Attachment 1: ETMY-ISCT_EISOL.jpg
ETMY-ISCT_EISOL.jpg
  9917   Tue May 6 17:58:44 2014 ericqUpdateLSCfarther into CM

 I took a look at the MC OLTF and AO path TFs with the fast agilent analyzer. 

I played with the relative gain of the EOM and PZT, but couldn't really change the MC OLTF shape much without making the PC Drive RMS angry. 

However, it turns out we have plenty of phase headroom to up the MC UGF from ~100kHz to ~180, with about 40 degrees of phase margin and ~7dB of gain margin. As I write this, PC drive RMS is around 1.1, and FSS Fast at 5.6, so I think the extra gain is fine for now. 

This pushes up and smoothens out the gain peaking in the AO path; see this figure:

AOTFs.pdf

(why does ELOG hate my python plots?! argggg)

Rana's rule of thumb was "We need at least +3dB MC loop gain at our CM servo UGF," so it looks like high tens of kHz bandwidth may be doable from the AO standpoint.

RXA: No, no, no, no, no, noooo. Rana said we need a gain of 3-10 at the CM UGF, not +3 dB.

  9916   Tue May 6 10:31:58 2014 jamieUpdateCDSc1ioo dolphin fiber

Quote:

I put label  at the dolphin fiber end at 1X2 today.   After this I had to reset it, but it failed.

 If by "fail" you're talking about the c1oaf model being off-line, I did that yesterday (see log 9910).  That probably has nothing to do with whatever you did today, Steve.

  9915   Tue May 6 10:22:28 2014 steveUpdateCDSc1ioo dolphin fiber

Quote:

Steve and I nicely routed the dolphin fiber from c1ioo in the 1X2 rack to the dolphin switch in the 1X4 rack.  I shutdown c1ioo before removing the fiber, but still all the dolphin connected models crashed.  After the fiber was run, I brought back c1ioo and restarted all wedged models.  Everything is green again:

green.png

 I put label  at the dolphin fiber end at 1X2 today.   After this I had to reset it, but it failed.

Attachment 1: dolphin1x2.png
dolphin1x2.png
  9914   Tue May 6 09:46:50 2014 steveUpdatePEMcleaning day

 Keven and Steve,

We cleaned around the Vertex chambers, PSL and MC2 floor areas.

 

Attachment 1: 120d.png
120d.png
  9913   Tue May 6 03:17:15 2014 ranaUpdateLSCfarther into CM

Yes, we still need to do these things, day team. Please tune up the MC loop first, before anything else.

Quote:

Next steps:

  • Check CM board boosts turn on politely (Transients, TFs)
  • Use fast spectrum analyzer to check MC loop gain out to a few MHz. (The bump in the tens of kHz should be fixed / moved higher)
  • Think about noise performance of, say, REFLDC, ASDC, RF AS signals, etc. in the PRY case, figure out which one to use first. 
  • We may want to first focus on directly locking the arm on an RF signal, figure out gains etc. and then figure out how to do DC->RF handoff nicely, or if high bandwidth DC signal control is even feasible.  

  9912   Tue May 6 02:48:50 2014 JenneUpdateLSCAO path engaged with AS55 as error signal for Yarm locking

[Rana, Jenne]

This evening, we were able to lock the Yarm through the common mode board, using AS55 as our error signal.  Our final UGF is about 5kHz, with 60 degrees of phase margin.

Before dinner, Rana switched the input of the CM board's REFL1 input to be AS55I rather than POY11Q, in the hopes that it would have better SNR.  Demod phase of AS55 was measured to be 14 deg for optimum Yarm->I-phase and has been set to 0 degrees.  Since the POY demod phase had been 90 degrees, which puts in a minus sign, and now we're using 0 deg which doesn't have a minus sign, we're using the plus (instead of minus) polarity of the CM board.

We re-allocated gains to help lower the overall noise by moving 15dB from the CM board AO gain slider to the MC IN2 gain slider, so we weren't attenuating signals.

We see, by taking loop measurements even before engaging the AO path (so, just the digital loop portion) that we've gained something like 20 degrees of phase margin!  We think that about 5 degrees is some LSC loop re-shaping of the boost filter.  We weren't sure why there was a hump of extra gain in the boost filter, so we've created a new (FM8) boost filter which is just a usual resonant gain:  resgain(16.5,7,50)

The cm_down and cm_step scripts in ..../scripts/PRFPMI/ were modified to reflect the settings below, and their current states are included in the tarball attached.

Also, throughout our endeavors this evening, the PC fast rms has stayed nice and low, so we don't suspect any EOM saturation issues.


Now our Yarm digital servo has a gain of -0.0013, with FMs 2, 4, 5, 7, 8 engaged (2, 7, 8 are triggered). 

Our final CM board settings are: 

REFL1 gain = +22dB

offset = -2.898V

Boost = enable

Super Boost = 0

option = disable

1.6k:79 coupled cavity compensator = enabled

polarity = plus

option = disable

AO gain = 15dB

limiter = enable

MC board:  IN1 gain = 18dB, IN2 gain = 0dB.


Here is a measurement of the Common Mode MCL/AO crossover.  The purple/orange trace here is after/before the boost was engaged.

out.pdf

We also have a measurement of the total loop gain, measured with the SR785.  The parameter file, as well as the python script to get the data, are in the tarball attached.  Noteably, the excitation amplitude was 500mV, whereas Q and Rana yesterday were using 5 or 8 mV.  We aren't sure why the big change was necessary to get a reasonable measurement out.  This measurement is with the boost enabled.

TF3_5May2014_BoostON_UGF5kHz.png

Finally, here is a measurement of the MC error point spectra, with the CM boost on, after we reallocated the gains.  There's a giant bump at several tens of kHz.  We need to actually go out with the fast analyzer and tune up the MC loop.

CM_TP2A_140506_boostON_realloc.png

Attachment 2: zipped.tgz
  9911   Mon May 5 19:51:56 2014 jamieUpdateCDSc1oaf model broken because of broken BLRMS block

I finally tracked down the problem with the c1oaf model to the BLRMS part:

/opt/rtcds/userapps/release/cds/common/models/BLRMS.mdl

blrms-hot-mess.pngsddefault.jpg

Note that this is pulling from a cds/common location, so presumably this is a part that's also being used at the sites.

Either there was an svn up that pulled in something new and broken, or the local version is broken, or who knows what.

We'll have to figure how what's going on here, but in the mean time, as I already mentioned, I'm leaving the c1oaf model off for now.

 RXA: also...we updated Ottavia to Ubuntu 12 LTS...but now it has no working network connection. Needs help.  (which of course has nothing whatsoever to do with this point )

  9910   Mon May 5 19:34:54 2014 jamieUpdateCDSc1ioo/c1ioo control output IPCs changed to PCIE Dolphin

Now the c1ioo in on the Dolphin network, I changed the c1ioo MC{1,2,3}_{PIT,YAW} and MC{L,F} outputs to go out over the Dolphin network rather than the old RFM network.

Two models, c1mcs and c1oaf, are ultimately the consumers of these outputs.  Now they are picking up the new PCIE IPC channels directly, rather than from any sort of RFM/PCIE proxy hops.  This should improve the phase for these channels a bit, as well as reduce complexity and clutter.  More stuff was removed from c1rfm as well, moving us to the goal of getting rid of that model entirely.

c1ioo, c1mcs, and c1rfm were all rebuild/installed/restarted, and all came back fine.  The mode cleaner relocked once we reenabled the autolocker.

c1oaf, on the other hand, is not building.  It's not building even before the changes I attempted, though.  I tried reverting c1oaf back to what is in the SVN (which also corresponds to what is currently running) and it doesn't compile either:

controls@c1lsc ~ 2$ rtcds build c1oaf
buildd: /opt/rtcds/caltech/c1/rtbuild
### building c1oaf...
Cleaning c1oaf...
Done
Parsing the model c1oaf...
YARM_BLRMS_SEIS_CLASS TP
YARM_BLRMS_SEIS_CLASS_EQ TP
YARM_BLRMS_SEIS_CLASS_QUIET TP
YARM_BLRMS_SEIS_CLASS_TRUCK TP
YARM_BLRMS_S_CLASS EpicsOut
YARM_BLRMS_S_CLASS_EQ EpicsOut
YARM_BLRMS_S_CLASS_QUIET EpicsOut
YARM_BLRMS_S_CLASS_TRUCK EpicsOut
YARM_BLRMS_classify_seismic FunctionCall
Please check the model for missing links around these parts.
make[1]: *** [c1oaf] Error 1
make: *** [c1oaf] Error 1
controls@c1lsc ~ 2$ 

I've been trying to debug it but have had no success.  For the time being I'm shutting off the c1oaf model, since it's now looking for bogus signals on RFM, until we can figure out what's wrong with it. 

Attachment 1: ioo-ipc.png
ioo-ipc.png
  9909   Mon May 5 19:26:43 2014 JenneUpdateLSCOverride ability for whitening triggering

 Today I finally implemented a feature in the whitening triggering that I should have a long time ago:  an override button.

Now, on each RFPD's phase rotation screen, there is a button to either allow triggering for that PD (both quads) or to be in manual mode.  

If you are allowing triggering, they will behave as they have for the last ~year.....if any degree of freedom is using either quadrant of that PD, and that degree of freedom is triggered, then engage analog whitening and digital de-whitening.

If you chose manual mode, then you can engage or disengage the whitening as you please.  (The analog whitening and digital de-whitening are still tied together).

  9908   Sun May 4 22:28:54 2014 ericqUpdateLSCfarther into CM

 [Rana, ericq]

Today, we got a ~2kHz bandwidth lock of the YARM with the AO path. We weren't able to turn any boosts on, due to POY noise. 

Rana and Koji have written scripts (/scripts/PRFPMI/cm_step and cm_down) that work very reliably. 

Here is an OLTF. (Violin filter was off, the crap around 600Hz goes away with them on)

 OLTF.png

My MATLAB modeling was useful is predicting the features of the loop shape, and the dependence on AO gain/crossover. Still, I need to check it out, because there is nonzero discrepancy between reality and my model (this may be hiding in the non flat MC AO response, i.e. the bump at ~35kHz. Alternatively, the crossover frequency is a free parameter...)

In any case, we have confidence that the CM board is mostly working predictably. We presume that our current obstacle is the very noisy nature of POY, and thus it's not worth spending more time in this configuration. 

Upcoming plans:

  • Use the CM board to control the Y arm coupled with the PRM. ("PRY"?)
  • Determine the game plane for high BW control of CARM. 

Next steps:

  • Check CM board boosts turn on politely (Transients, TFs)
  • Use fast spectrum analyzer to check MC loop gain out to a few MHz. (The bump in the tens of kHz should be fixed / moved higher)
  • Think about noise performance of, say, REFLDC, ASDC, RF AS signals, etc. in the PRY case, figure out which one to use first. 
  • We may want to first focus on directly locking the arm on an RF signal, figure out gains etc. and then figure out how to do DC->RF handoff nicely, or if high bandwidth DC signal control is even feasible. 

RXA: we should also use AS45 instead of POY11. It has better SNR and I think our whole problem is too little light on POY.

  9907   Sun May 4 14:20:04 2014 ericqUpdateIOOPMC relocked

The PMC has been unlocked for ~23 hours. FSS slow was at ~-1.5 V. I zeroed it, and relocked the PMC, transmission is ~0.81V. MC with WFS came back fine.

  9906   Fri May 2 19:03:13 2014 JenneUpdateLSCALS out of loop spectra

I have taken out of loop spectra for both arms, by looking at POX/POY while the arms were controlled with ALS.

To do this, I put the POY11 Q signal directly to the whitening board, which meant that I removed the cable coming from the common mode board.  (Now that we're doing CM stuff again, I have put it back, so POY is still in the slightly weird "going through the CM slow path" situation). 

For the locking, both arms had FMs 1, 2, 3, 5, 6 engaged.  Yarm had a gain of +17, Xarm had a gain of -17. 

Y beatnote was 98.6MHz with a peak height of -22 dBm.  X beatnote was 45.0MHz with a peak height -11 dBm.

I drove ITMY at 503.1 Hz with 100 counts.  I drove ITMX at 521.1 Hz with 25 cnt. 

Koji helped me match up the peak heights between the FINE_PHASE_OUT_HZ calibrated signals and the PDH signals. 

The out of loop noise is definitely below 1kHz rms now, which is better than it was!  Hooray!

ALS_OutOfLoop2_2May2014.pdf

  9905   Fri May 2 14:31:26 2014 KojiUpdateLSCALS Y beat setup aligned

Please check the X&Y ALS out-of-loop stability. Use fine resolution (BW0.01). Calibrate the POX/POY in Hz.

  9904   Fri May 2 13:03:30 2014 JenneUpdateLSCALS Y beat setup aligned

I touched up the alignment of the Ygreen on the PSL table.

  9903   Fri May 2 11:14:47 2014 jamieUpdateCDSc1ioo dolphin fiber nicely routed

Quote:

This C1IOO business seems to be wiping out the MC2_TRANS QPD servo settings each day.   What kind of BURT is being done to recover our settings after each of these activities?

(also we had to do mxstream restart on c1sus twice so far tonight -- not unusual, just keeping track)

I don't see how the work I did would affect this stuff, but I'll look into it.  I didn't touch the MC2 trans QPD signals.  Also nothing I did has anything to do with BURT.  I didn't change any channels, I only swapped out the IPCs.

  9902   Fri May 2 10:38:29 2014 steveUpdatePSLthin window AR measured

Quote:

CVI broadband AR coating was measured at the PSL-enclosure table around 9-10am today. The 2W Innolight first PBS  S polarization beam was used with an other 1/2 wave plate and PBS.

W2-PW1-1004-C-633-1064-0   This 0.045" thick window has 0.7- 0.8 % reflected beam on each sides at 5 degrees of incidence, P polarization.

The specification is  R avg <0.5 % per surface at 0 degree

Rana wants The device would be useless with such a high R, but R 0.1% is OK so I will get V coating.

 CVI V-AR coating at 1064 nm, 0 degree,  catalog item is R< 0.25% on each sides,

 R <0.1 % is custom at much higher prices.

This custom order should go with other  orders that has similar need.

From CVI: 5-6-2014

I checked the trace info on the W2-PW1-1004-C-633-1064-0, BBAR coated window that you received.  It is side 1, 0.42%R & side 2, 0.53%R @ 1064nm.  And with the shift, I’m not too surprised you ended up with 0.7%.  A V coat would start with <0.25% (and more typically coming in at ~0.1%) per surface.  As far as stock options, I have a 1”dia x 4mmT, fused silica window that is recorded as side 1, R=0.09 and Side 2, R=0.08% @ 1064.  Is this too think or will it work for you?

 

 

  9901   Fri May 2 08:38:07 2014 SteveUpdateIOOthis typical morning

 

 c1sus needed reset.

Attachment 1: 2dTrend.png
2dTrend.png
  9900   Fri May 2 08:15:54 2014 steveUpdatesafetysafety audit 2014

 

 Late adition: CHECK all viewport covers.

A, transparent Lexan sheet is protecting glass windows in a horizontal position

B, metal housing protection is required on each viewport except signal ports

C, signal ports should be shielded by optical table enclosure

 

We have to cover this window-camera with implosion proof cover or just remove it and blank it.

Question number 2: Do our vertically positioned windows with flip able covers require protective lexan ? NO 5-5-2014

Attachment 1: solidCoverRequired.jpg
solidCoverRequired.jpg
  9899   Fri May 2 03:51:29 2014 ranaUpdateLSCfarther into CM

Rana, Q

After some more matlab loopology (see Qlog), we turned on the AO path successfully. The key was to turn on the 300:80 filter in the MCL path so that it could cross stably with the AO. Then we ramp up the AO gain via the newly AC coupled AO path into the MC servo board.

The POY11 signal looks nice and smooth. For the final smoothness after the overall common gain is ramped up, I turned on a FM7 pole at 300 Hz so that the MC path would keep falling like 1/f^2 and not interfere with the AO path around 1 kHz.

There's not enough gain yet to be able to turn on the Boost. PCDRIVE is ~3 V. Earlier tonight we were seeing the EOM saturation effect maybe, but we re-allocated the gain more to the front end and its all fine now. I think we can get another ~10-15 dB of gain by using the POY whitening gain slider + the CM AO slider. Then we can get the Boost on and take some TFs with the SR785 (as long GPIB allows).

Good Settings:

CM REFL1 = +31 dB, AO = +16 dB, MC IN2 = +16 dB. SUS-MC2_LSC = FM6, FM& ON

 

** Everything has been pretty stable tonight except some occasional MC/EOM locking oscillations. This means that its been easy to keep trying some different CM steps since the Y-Arm relocks using MCL within seconds.

Attachment 1: MCkicked.png
MCkicked.png
  9898   Fri May 2 02:22:56 2014 rana, QUpdateIOOMC alignments

We were having unstable MC locking so we did some physical alignment touch up. Use MC suspension bias to have good MC alignment. Unlock MC and align DC beams to center on WFS. Re-lock and things are now stable.

Someone had been feeding bad info to Eric about MC alignments; we do not center the MC WFS beams with the MC locked.

However, in any case, today the MC2-TRANS servo was not being good so I turned it off. We need the real matrix measurement to bring it back.

  9897   Fri May 2 01:58:52 2014 ranaConfigurationComputer Scripts / ProgramsupdateDB configured to index NFS during CRON daily

I wanted to use locate to find some files today, but found that it doesn't index any of our NFS mounted files (i.e. the whole /cvs/cds/ and /opt/rtcds/ ).

So I went into /etc/updatedb.conf and edited it like so, so that it no longer ignores NFS type file systems:

controls@rossa:/etc 0$ diff updatedb.conf updatedb.conf~
4c4
< PRUNEFS="nfs4 rpc_pipefs afs binfmt_misc proc smbfs autofs iso9660 ncpfs coda devpts ftpfs devfs mfs shfs sysfs cifs lustre_lite tmpfs usbfs udf fuse.glusterfs fuse.sshfs ecryptfs fusesmb devtmpfs"
---
> PRUNEFS="NFS nfs nfs4 rpc_pipefs afs binfmt_misc proc smbfs autofs iso9660 ncpfs coda devpts ftpfs devfs mfs shfs sysfs cifs lustre_lite tmpfs usbfs udf fuse.glusterfs fuse.sshfs ecryptfs fusesmb devtmpfs"

The CRON.daily on ROSSA only should run each morning at 6:25 (under the ionice class 3 protocol as usual). If this seems OK after a week or so, we can do the same for the other CDS workstations (remembering to adjust their cron.daily times so that not every one hammers the NFS at the same time).

  9896   Fri May 2 01:01:28 2014 ranaUpdateCDSc1ioo dolphin fiber nicely routed

This C1IOO business seems to be wiping out the MC2_TRANS QPD servo settings each day.   What kind of BURT is being done to recover our settings after each of these activities?

(also we had to do mxstream restart on c1sus twice so far tonight -- not unusual, just keeping track)

  9895   Thu May 1 17:14:36 2014 steveUpdatePSLthin window AR measured

CVI broadband AR coating was measured at the PSL-enclosure table around 9-10am today. The 2W Innolight first PBS  S polarization beam was used with an other 1/2 wave plate and PBS.

W2-PW1-1004-C-633-1064-0   This 0.045" thick window has 0.7- 0.8 % reflected beam on each sides at 5 degrees of incidence, P polarization.

The specification is  R avg <0.5 % per surface at 0 degree

Rana wants The device would be useless with such a high R, but R 0.1% is OK so I will get V coating.

  9894   Thu May 1 17:00:05 2014 ranaUpdateLSCYarm locking with CM board

 I think that's about halfway there. Since this needs to be a precise comparison, we cannot use so many approximations.

We've got to include the digital AA and AI filters as well as the true, measured, time delay in the system. Also the measured/fitted TF of the CM board with the 79:1.6k filter engaged. We want an overall phase accuracy between Jenne's measured TF from last night and this model (i.e. on the same plot with the residual plotted).

Is there a cavity pole in the model? Should be at ~1.6 kHz.

  9893   Thu May 1 16:41:35 2014 ericqUpdateLSCYarm locking with CM board

 (Edited this post; Forgot to account for the FMs other than 4 and 5... it now agrees better!)

I did some quick MATLAB simulation of the relevant loops to try and understand what was going on. I put the digital UGF around 200Hz, and then brought in the AO path with both signs. 

In these plots, blue is digital only, green is AO+digital with the crossover happening at the UGF, and red is the AO gain set to five times of what it was in the green curve. 

 AOsignsSame.pdfAOsignsOpposite.pdf

Based on the phase curves in the loop measurements, I would be inclined to say the pink -AO case corresponds to the opposite sign plot, and the +AO case to the same sign plot. 

This correspondence also holds for the appearance of the peaks in the noise curves, the Opposite sign case has a dip in loop gain at ~50Hz (pink curve, -AO), same sign around ~30Hz (brown curve, +AO). 

However, both of these look like they become unstable at some point in the transition! This agrees with our experience last night...

I'll fiddle around and try to come up with some compensating digital filter that will make the Opposite sign scenario work. 

The MATLAB code I used to make these plots is attached. 

Attachment 3: loopModeling.m
clear all

cycleT = 60e-6;

% AI, AA shapes from http://nodus.ligo.caltech.edu:8080/40m/8555
[z,p,k] = ellip(4,4,60,2*pi*7570,'s');
AI = zpk(z,p,k*10^(4/20)) * zpk([],-2*pi*13e3,2*pi*13e3);
AI.OutputDelay = 1*cycleT;

[z,p,k] = ellip(8,0.001,80,2*pi*7570,'s');
... 58 more lines ...
  9892   Thu May 1 14:45:44 2014 JenneUpdateLSCMC board back in

Quote:

Quote:

To fix this, I am going to put a 6.8uF cap in series with R30 in the MC board, which is part of the crossbar switch where the IN1 and IN2 get summed.  This should AC-couple the output of the IN2 slider before the summing node.

 MC board is out, so don't be surprised that the MC isn't locking.

 MC board is back in place, MC is locked.

If I disable all of the AO path bits of the CM servo (disable switch, and also gain slider to -32dB), and then move the MC IN2 slider around, the MC does not get an offset anymore (as seen by reduced transmission and increased reflected power), so I think the DC coupling is working.  I do lose lock of the MC if the slider goes above ~22 dB in this situation, but I don't see any effect before then, whereas we were able to see a steady increase in the reflected power as we moved the slider around last night.  So, it seems like things are good with the DC coupling of the IN2 slider.

Here are some photos from before I modified the board (front, back, and zoom of the area I was working in):

IMG_1394.JPG

IMG_1395.JPG

IMG_1398.JPG

And here is my modification, putting the 6.8uF cap in series with (a new) 2k thin film resistor, in the spot for R30:

IMG_1402.JPG

The board is https://dcc.ligo.org/DocDB/0004/D040180/001/D040180-C.pdf

[Edit, 20140721: It looks like this is actually D040180 rev B, not rev C. —Evan]

  9891   Thu May 1 13:03:34 2014 JenneUpdateLSCMC board pulled for AC coupling

Quote:

To fix this, I am going to put a 6.8uF cap in series with R30 in the MC board, which is part of the crossbar switch where the IN1 and IN2 get summed.  This should AC-couple the output of the IN2 slider before the summing node.

 MC board is out, so don't be surprised that the MC isn't locking.

  9890   Thu May 1 10:23:42 2014 jamieUpdateCDSc1ioo dolphin fiber nicely routed

Steve and I nicely routed the dolphin fiber from c1ioo in the 1X2 rack to the dolphin switch in the 1X4 rack.  I shutdown c1ioo before removing the fiber, but still all the dolphin connected models crashed.  After the fiber was run, I brought back c1ioo and restarted all wedged models.  Everything is green again:

green.png

  9889   Thu May 1 03:23:07 2014 ericqUpdateLSCYarm locking with CM board

Quote:

POY is going from its demod board to the CM board, and then the slow output of that is going to the POY channel of the whitening, and then on to the ADC.  So, with no AO path engaged, this is basically like regular Yarm locking.  

Just to be clear, the normal POY signals are not currently present, so the restore POY script will not result in the arm locking. POY11_I is turned off, POY11_Q is the output of the CM board, which can be used to lock the arm, as we did tonight. 

The POY digital demos angle went -56 -> 90, to get all of POY11_Q_IN1 to POY11_I_ERR

Miscellaneous things:

  • Right now, the cable from CM board ->MC board is a BNC. There appeared to be a differential 2-pin lemo hanging around for this purpose, but it didn't seem to be transmitting the signal. However, we will want something better than a BNC to keep this signal clean. 
  • I took SR785 TFs of the CM board from IN to the slow and fast outs. They looked reasonable, and will be posted in time. 
  • We enabled the 79:1.6k filter in the CM screen (though it is unclear if these are the actual values...), and put in its inverse in the digital path. I.e. we only want this shape in the AO path, to give it 1/f shape in the vicinity of the crossover. This is only necessary in the uncoupled cavity case. 
  9888   Thu May 1 03:15:03 2014 JenneUpdateLSCYarm locking with CM board

[Rana, EricQ, Jenne]

We locked the Yarm by using the CM board this evening. 

POY is going from its demod board to the CM board, and then the slow output of that is going to the POY channel of the whitening, and then on to the ADC.  So, with no AO path engaged, this is basically like regular Yarm locking. 

First of all, Den and Koji back in December were concerned that they were seeing some EOM saturation in the fast path, but we don't think that's an issue.  We looked at the FSS PCDRIVE while we increased the AO gain.  In fact, it looks like the offset is coming from the MC board's IN2 slider.  Even with no input on that slider, increasing its value puts an offset into the MC.  To fix this, I am going to put a 6.8uF cap in series with R30 in the MC board, which is part of the crossbar switch where the IN1 and IN2 get summed.  This should AC-couple the output of the IN2 slider before the summing node.

We aren't sure which sign to use for the AO path of the CM board...Eric is doing some modelling to see if he can figure it out.  He's going to try to see which spectra (below) his model matches.

For the spectra, we have a reference trace with no AO path, a trace with "Plus" polarity on the CM board which started to show a peak when the value of the MC IN2 slider was at about -6 dB, and a trace with "Minus" polarity, which started to show a peak when the value of the MC IN2 slider was at about -16 dB. 

Yarm_CMlocking_spectra_30Apr2014_copy.pdf

We took loop measurements for each of the Plus and Minus cases. Something that seems a little weird is how shallow of a slope we have in both cases near our UGF.

Yarm_CMlocking_TFs_30Apr2014_copy.pdf

 

  9887   Thu May 1 00:13:21 2014 KojiUpdateLSCALS X beat setup aligned

I saw big misalignment on the GTRX camera, I went to the PSL table and aligned the beat beams.

I disconnected the RF out of the X beat PD and  connect an oscilloscope.
The beat amplitude was 15mVpp at the beginning and is 60mVpp right now.
I checked the alignment on this RF PD and the DC PD as well as the spot on the CCD.

The RF cable was connected again.

Jenne and I ran the ALS and scanned the arm cavity. We had the impression that the noise level of the ALS improved,
but I don't have correctly calibrated measurement. Let's do it tomorrow in the day time.

The Yarm beat alignment look awful. We should align this too.

  9886   Wed Apr 30 21:57:07 2014 ranaSummaryIOOMC2_TRANS QPD Servo now on for real

dolphin.pngMC2_QPD_trend.png

During a lull in Koji vs. The Arm, I switched on the MC2_TRANSQPD feedback path to check out its UGF. In the past months, when its been on, it has had a gain of ~0.03 - 0.1.

Today, I found that with the gain turned up to 11, it has a ~1 minute step response time (as you see in the above Strip chart). So its had a UGF of ~2 hours or so during the times when we thought it might be doing bad or good or magic.

I leave it on now to see if it behaves well over the next days. Let's see if Steve thinks its good or not based on his trend monitoring.

** also touched up the PMC pointing (using the PMC REFL image / please never align the beam into the PMC without this camera image)

  9885   Wed Apr 30 21:31:25 2014 ranaHowToIOOMystery Alignment again

Looks like there was some mysterious MC alignment shift around 5:30 PM today, but no elog.....?? Now things are drifting much more than this morning or yesterday. Who did what and why???

I think I'll blame Jamie since he just got back and did some computer fiber voodoo today.

http://www.lawsome.net/no-throwing-rotten-tomatoes-a-repealed-kentucky-law/

  9884   Wed Apr 30 21:16:42 2014 ranaUpdateLSCMC2 Strad

bettsreplica.jpg

I found the YARM LSC feedback going to MC2 and the MC2 violin mode (at 644.69 Hz) rung up. The existing notch was just a second order Twin-T style notch (so not a good idea) and also not turned on, since it was in the FM4 spot of LSC-MC2 and the vio triggers are ganged between all mirrors and don't touch FM4.

I copied the PRM vio bandstop into FM2 of this bank, deleted the old notch, and tuned the bandstop frequencies a little to get the violin peak into one of the zeros of the elliptic bandstop. Attached are the Y-arm / MCF spectrum with the mode rung up as well as the new filter's TF compared with the old notch.

P.S. I installed http://en.wikipedia.org/wiki/Midnight_Commander on pianosa.

Attachment 2: MC_Y_vio.pdf
MC_Y_vio.pdf
Attachment 3: MC2_vio.pdf
MC2_vio.pdf
  9883   Wed Apr 30 18:06:06 2014 jamieUpdateCDSPOP QPD signals now on dolphin

The POP QPD X/Y/SUM signals, which are acquired in c1ioo, are now being broadcast over dolphin.  c1ass was modified to pick them up there as well:

c1ioo-POPQPD.pngc1ass-POPQPD.png

Here are the new IPC entries:

controls@fb ~ 0$ egrep -A5 'C1:IOO-POP' /opt/rtcds/caltech/c1/chans/ipc/C1.ipc
[C1:IOO-POP_QPD_SUM]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=116
desc=Automatically generated by feCodeGen.pl on 2014_Apr_30_17:33:22
--
[C1:IOO-POP_QPD_X]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=117
desc=Automatically generated by feCodeGen.pl on 2014_Apr_30_17:33:22
--
[C1:IOO-POP_QPD_Y]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=118
desc=Automatically generated by feCodeGen.pl on 2014_Apr_30_17:33:22
controls@fb ~ 0$ 

Both c1ioo and c1ass were rebuild/install/restarted, and everything came up fine.

The corresponding cruft was removed from c1rfm, which was also rebuild/installed/restarted.

  9882   Wed Apr 30 17:45:34 2014 jamieUpdateCDSc1ioo now on Dolphin network

For reference, here are the new IPC entries that were made for the ALS X/Y phase between c1als and c1lsc:

controls@fb ~ 0$ egrep -A5 'C1:ALS-(X|Y)_PHASE' /opt/rtcds/caltech/c1/chans/ipc/C1.ipc
[C1:ALS-Y_PHASE]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=114
desc=Automatically generated by feCodeGen.pl on 2014_Apr_17_14:27:41
--
[C1:ALS-X_PHASE]
ipcType=PCIE
ipcRate=16384
ipcHost=c1ioo
ipcNum=115
desc=Automatically generated by feCodeGen.pl on 2014_Apr_17_14:28:53
controls@fb ~ 0$ 

After all this IPC cleanup is done we should go through and clean out all the defunct entries from the C1.ipc file.

  9881   Wed Apr 30 17:07:19 2014 jamieUpdateCDSc1ioo now on Dolphin network

The c1ioo host is now fully on the dolphin network!

After the mx stream issue from two weeks ago was resolved and determined to not be due to the introduction of dolphin on c1ioo, I went ahead and re-installed the dolphin host adapter card on c1ioo.  The Dolphin network configurations changes I made during the first attempt (see previous log in thread) were still in place.  Once I rebooted the c1ioo machine, everything came up fine:

dolphin.png

We then tested the interface by making a cdsIPCx-PCIE connection between the c1ioo/c1als model and the c1lsc/c1lsc model for the ALS-X beat note fine phase signal.  We then locked both ALS X and Y, and compared the signals against the existing ALS-Y beat note phase connection that passes through c1sus/c1rfm via an RFM IPC:

The signal is perfectly coherent and we've gained ~25 degrees of phase at 1kHz.  EricQ calculates that the delay for this signal has changed from:

ALSXonDolphin.pdf

122 us -> 61 us 

I then went ahead and made the needed modifications for ALS-Y as well, and removed ALS->LSC stuff in the c1rfm model.

Next up: move the RFM card from the c1sus machine to the c1lsc machine, and eliminate c1sus/c1rfm model entirely.

  9880   Wed Apr 30 16:07:59 2014 manasaUpdateLSCALS X noise post servo modification

I. The Y arm stayed stable through last night and I have saved the arm lock settings to IFOconfigure.

II. ALS X arm noise measurements

I looked at the before and after noise of ALSX.

Settings:
Phase tracker gain = 85
Xarm servo gain = -17

The rms in loop noise has dropped from 3KHz to 500 Hz.

Attachment 1: Phase tracker OLTF

Attachment 2: Free running noise and in loop noise

Attachment 3: Out of loop noise (measured with arms locked using PDH for IR)

Attachment 4: ALS arm servo OLTF

xml data files can be found in /users/manasa/data/140430/

Attachment 1: ALSX_PToltf.jpg
ALSX_PToltf.jpg
Attachment 2: ALSX_FreeInloop.jpg
ALSX_FreeInloop.jpg
Attachment 3: ALSX_ool.jpg
ALSX_ool.jpg
Attachment 4: ALSX_OLTF.jpg
ALSX_OLTF.jpg
  9879   Wed Apr 30 14:21:50 2014 manasaUpdateCDSfb restarted

c1sus and c1isey were not talking to fb. The usual mxstream restart did not help.

Restarted fb

>>ssh fb

>>telnet fb 8087
shutdown

All lights on the FE status screen are green now.

Note that Steve did an mxstreamrestart earlier today because the same machines c1sus and c1isey were not talking to fb.

  9878   Wed Apr 30 11:16:41 2014 SteveUpdateIOOthis typical morning

c1sus and c1iscey were reset. The PMC needed to be locked. MC locked instantly.

The FSS ion pump power supply was turned on.

 

 

 

 

Attachment 1: 40days.png
40days.png
  9877   Wed Apr 30 00:40:55 2014 manasaConfigurationLSCY arm IR lock troubleshooting

[Koji, Manasa]

The Y arm locks stably for IR PDH now.  

The reason for ETMY getting kicked during lock acquisition was finally found to be related to the limiter value set in the Y arm servo. We reduced the limiter value unintentionally and found that the lock acquisition stayed smooth. The limiter value was stepped in 1000s from 7000 and eventually found that the ETMY suspension was kicked when we try to acquire lock with the limiter value was set at 11000. 

 

The limiter for X arm at 11000 is not causing any trouble at the moment.

In the process, we did a bunch of things through the evening to troubleshoot IR locking of the Y arm.

Earlier today running the IFO configure script did not restore the arms to lock and both the ETMs needed to be aligned to lock the arms. The arms stayed locked for 15 minutes and the Y arm lost lock eventually leaving the ETMY in a misaligned state. 

The state of the Y arm was similar to what Jenne has explained in ELOG where the ETMY was kicked during lock acquisition and would move to a misaligned state.

To trouble shoot, there were several things that were done. A few of them might not have any direct correlation to the locking issue but could just be a coincidence.  

1.  The trigger time for the filters in the arm filter modules were set such that they switch on after the SUS violin filters. Arm FM trigger time = 3 s (previously set at 0.1s) and SUS violin trigger time = 1s. This reduced the number of lock loss events.

2. There was some drop in transmission when the bounce filter of Y arm (FM6) turned ON. This was fixed by changing its ramp time (initially set at 1s). The filter has been modified to turn on immediately upon arm lock acquisition before the other triggered filters in the filter module turn on.

3. The QPD and SUS signal cables running to the rack were checked to be intact. Koji found some of them to be loose. But this had no evident correlation with the arm locking problem.

4.The oplev and PD alignment was checked at the Y end. The high gain trans PD for Y arm was checked for good alignment by looking at TRY. It was found that the EXIT light at the Y end is injecting some noise to the transmission PD. 

5. The ETMY was given offsets in PIT, YAW and POS and the OSEM sensor values were checked to see if the suspension is behaving well. It was behaving well.

  9876   Tue Apr 29 16:42:29 2014 SteveUpdateVACTP2 drypump replaced

Quote:

 

 TP2's fore line - dry pump replaced at performance level 600 mTorr after 10,377 hrs of continuous operation.

Where are the foreline pressure gauges? These values are not on the vac.medm screen.

The new tip seal dry pump lowered the small turbo foreline pressure 10x

TP2fl after 2 day of pumping 65mTorr

 TP2 dry pump replaced at fore pump pressure 1 Torr,  TP2 50K_rpm 0.34A

 Top seal life 6,362 hrs

 New seal performance at 1 hr  36 mTorr, 

 Maglev at 560 Hz, cc1 6e-6 Torr

 

Attachment 1: dryforepumpreplaced.png
dryforepumpreplaced.png
  9875   Tue Apr 29 10:01:25 2014 KojiConfigurationGeneralnetgpibdata is working again now

I've moved the WB network analyzer to the OMC lab. The 40m network analyzer is not in service for the MC monitoring.
I setup the configuration so that the same command gives us the same spectrum measurement.

ELOG V3.1.3-